OpenStudy (anonymous):

Tessa invested $143 dollars into an account that earned interest at an annual rate of 8%. How much money will she have in the account after 5 years?

OpenStudy (johnweldon1993):

M = Mi (1+r)^t M=money Mi = money initially r = rate (in decimal form) t = time in years so M = 143 (1 + .08)^5 M = 143(1.08)^5 is?

OpenStudy (johnweldon1993):

i take it you got it? :)